The Top Quark Search

Finding Truth by Searching for Beauty

1. The Current Paradigm—The Standard Model 2. How We “Discover” Quarks

3. Why Must the Top Quark Exist?

4. Searching for Top at CDF

5. What to Look for in the Future

Pekka Sinervo University of Toronto Sep 22, 1993

Standard Model of Elementary Particles

Particle physics before ∼ 1960 had established

• Elementary particles appeared to be the basic constituents

of universe electron (e±) proton (p) neutron (n)

• Four fundamental interactions or forces:

1. Strong Force—bound nucleons (p, n) together

2. Electromagnetism—bound e− to nuclei

3. Weak Force—responsible for neutron β decay

4. Gravity—most obvious, but also the weakest

• Host of other “elementary particles”

– neutrinos (νe, νμ) (“...who ordered that?”)

– muon (μ)

– lots of other hadrons (π∘, K±, Δ’s,...) Seemed to be spare wheels....

Over the last 3 decades, pieces have come together:

• Hadrons are composite objects made up of quarks

• 3 “generations” of quarks and leptons

( \begin{matrix}u \\ d \end{matrix} ) ( \begin{matrix}c \\ s \end{matrix} ) ( \begin{matrix}t \\ b \end{matrix} )
( \begin{matrix}e \\ νe \end{matrix} ) ( \begin{matrix}μ \\ νμ \end{matrix} ) ( \begin{matrix}τ \\ ντ \end{matrix} )

⇔ ( \begin{matrix}Iz = +1/2 \\ Iz = −1/2 \end{matrix} )

• Three of the four forces can be described by relativistic quantum field theories:

– “bosons” mediate each force

EM  γ
Weak  W+ , W− , Z∘
Strong gluons

What’s Missing?

This picture is enormously successful

• Describes detailed behaviour of W and Z bosons

• Accurate description of Electroweak force

• Apparently QCD also well described

There are two missing pieces

1. The Higgs particle

2. The Top Quark (the 6th quark—call it what you want)

Why the Higgs?

The Higgs is what “unifies” the EM and Weak forces

• γ, W ± and Z∘ are part of one “gauge” group

• γ mediates the EM force – massless, giving us inverse-square law force

• W ± and Z∘ bosons mediate the weak force – masses come from interaction with a Higgs particle – gives uniquely “weak” nature to the force

The Higgs has not been found

• Searches for it exclude Higgs with mass < 60 GeV/c2 • Rationale for next generation of accelerators

Why do We Need a Top Quark?

1 Bottom quark appears to be member of an SU(2) doublet

• See this in scattering process e+ e− → bb̄

• Models with bottom as an SU(2) singlet predict flavour-changing neutral currents 2 EWK measurements imply existence of a “heavy” fermion

that couples to W and Z bosons

sin2 θW ≡ 1 − (\frac{MW}{MZ} )2 = \frac{πα}{√2GμMW 2}[\frac{1}{1 − Δr(Mtop, MH)} ]

where:

• MW and MZ are W and Z boson masses

• Gμ is μ decay constant

• Δr is “radiative” corrections

How Does One Find Quarks?

Quarks have unique properties

• come only in pairs (qq̄) or triplets (qqq)

• QCD predicts that we cannot find a free quark • come in three “colours”

• have different masses – Mu ∼ Md ∼ 150 − 200 MeV/c2 – Ms ∼ 500 MeV/c2

– Mc ∼ 1300 MeV/c2

– Mb ∼ 5000 MeV/c2

• can decay via the EWK force

– s → W −d

– c → W + s

– b → W −s

Up, Down and Strange

Existence of up, down and strange quarks inferred from

• baryon spectrscopy

• properties of elementary particles (n, p, π, ...)

• probes of interior of proton

First evidence for proton structure came from SLAC

• high energy electrons were scattered off a H2 target • rate of scatters at very large angles quite high

• fits the “parton” model quite well

Charm and Bottom

Charm discovered in e^{+} e^{−} annihilations

• experiments at SLAC and BNL discovered J/ψ resonances • consistent with cc̄ bound states

• Evidence also from study of the total e+ e− cross section

R = \frac{σ(e+ e− → anything )}{σ(e+ e− → μ+ μ−)}

Bottom discovered also as a bound state

• Υ resonances with mass of ∼ 10 GeV/c2

Top Search Techniques

First have to create top quarks

1. Pair production in e+ e− → tt

2. Pair production in pp → tt

The standard model top quark decays:

• semi-leptonically:

t → W + b
\underrightarrow{|} \{ \begin{matrix}e+ νe & (1/9) \\ μ+ νμ & (1/9) \\ τ+ ντ & (1/9) \end{matrix}

• hadronically:

t → W + b
\underrightarrow{|} \{ \begin{matrix}cs̄ & (1/3) \\ ud̄ & (1/3) \end{matrix}

Top events are characterized by an unusual number of energetic partons and are very “spherical”.

• At e+ e− colliders, spherical topology sufficient to tag top

• At hadron colliders, necessary to search for specific decays

– One t decaying semi-leptonically

– Both t’s decaying semi-leptonically

DELPHI collaboration (at LEP) searched for top by

• Selecting spherical events (Thrust < 0.9 and “minor”> 0.2)

• Observed 48 events and expected 59 ± 10 background

⇒ excludes Mtop < 44.5 GeV/c2 (95%CL) PL 242B, 536, 1990

Summary of e^{+} e^{−} top quark searches

• All the searches are insensitive to the detailed top decay modes

• Exclude a top quark with:

Mtop < 29.9 GeV/c2  (Topaz )
Mtop < 29.9 GeV/c2  (Amy )
Mtop < 30.2 GeV/c2  (Venus )
Mtop < 40.7 GeV/c2  (Mark II)
Mtop < 44.5 GeV/c2  (OPAL )
Mtop < 45.8 GeV/c2  (Aleph )
Mtop < 44.5 GeV/c2  (Delphi )

How Do We Really Find Top?

The “best” way of producing top quarks is in proton-antiproton collisions

• high energy-collisions can produce heavy particles • Can achieve high interaction rates

L ∼ 7 × 1030 cm−2s−1

• Cross section for top production is fairly large

Fermilab Tevatron collidor is highest energy p p collider

• 6 bunches of protons and 6 bunches of antiprotons

• Circulate in 6 km storage ring—6 interaction regions

• Two interaction regions are instrumented 1. CDF detector

2. D0 detector ⇒ p p interaction rate of ∼ 50 kHz

Tough part is that top quarks are produced rarely

• Many other processes also occur

• top quark is produced every 1 in 5 × 109 collisions

The CDF Experiment

Study of p p interactions at √s = 1.8 TeV

• 34 institutions

• 400+ physicists

The IPP Group consists of

McGill University  University of Toronto
Faculty  Ken Ragan  Pekka Sinervo
RA  Kurt Biery  Andreas Hölscher

Klaus Strahl

Grad Students Yunling Ye George Sganos
Kostas Kordas  Andreas Warburton
Peter Musgrave  Hyunsoo Kim

Wendy Taylor

Experiment just completed data collection for Run IA

• Run IA (9/92 through 5/93) 22 pb−1

• Run IB (11/93 through 6/94) ⇒∼ 75 pb−1

The CDF Detector

• general-purpose, 5000 ton detector

Detector consists of:

– Calorimeter to measure the energies of the particles – A magnetic spectrometer that measures charged

particle momenta

– A sophisticated real-time triggering system that records on average about 1 in every 50,000 collisions

Top Quark Search

CDF has been searching for top since 1987

• Current best preliminary limit is CDF dilepton analysis ⇒ 40 < Mtop < 113 GeV/c2 (95% CL)

Idea is to look for events where both top quarks decay semi-leptonically

• Will have two energetic leptons (e±’s or μ±’s)

• Few background sources

Observe two new events with these cuts

• Physics backgrounds are small (0.7 ± 0.2 events) – Z → ττ, bb̄ production

• Fake backgrounds is estimated to be 0.4 ± 0.1 events

To set limit

• Assume these two events arise from top production • Include it with 1988-89 eμ candidate

Exclude 28 < Mtop < 113 GeV/c2 (95%CL)

Lepton + Jets Search

Idea is to look for decays where

Select lepton + jet candidates by requiring

• Clean, isolated lepton candidate with PT > 20 GeV/c

• E̸T > 25 GeV to tag neutrino

• Require jets to have observed ET > 15 GeV

• Reject Z∘ and conversion backgrounds

Resulting sample is dominated by W +jets production

• Fake lepton backgrounds are < 10 %

• Most of the top signal would appear as 3 and 4 jet events

Get rid of W + jets background by requiring “beauty”

• Use two methods:

– b quarks have “long” lifetimes (cτ ∼ 10−12 s) ⇒ search for jets with long-lived particle

– b quarks decay semileptonically also ⇒ search for jets with leptons

Studied algorithms with “control” samples

• QCD jet sample (small inherent B content)

• Inclusive lepton samples (enriched B sample)

Results of tagging studies

• Tagging efficiencies are ∼ 0.3/event

• “Fake” Tags have small rate (∼ 1 % of events)

• Largest source of background from W + bb̄

Define “signal sample” as events with ≥ 3 jets

• Tag three events (all three SVX algorithms tag these!)

• Estimate fake tag rate at < ∼ 1 event

Can compare observed tag rate with expectation for top production and backgrounds

• Have not yet set limit given small number of signal events – Collaboration is still “mulling” it over

• Data does tend to favour a heavier top mass?

Future Collider Runs and CDF Upgrades

Fermilab Collider will begin Run IB in November 1993

• Run for ∼ 12 months

• New LINAC injector

• Double maximum luminosity to 1.5 × 1031 cm−2s−1

• Goal is for CDF/D0 to collect ∼ 75 pb−1

CDF Upgrades for Run IB

• Replace downstream end of DAQ system

• SVX′ detector (complete replacement for SVX)

• Trigger improvements and L3 upgrade (×2)

Conclusions

• Search for Top Continues

• Standard Model top mass > 113 GeV/c^{2} • Lepton + Jets analysis not quite complete

→ Wait for Run IB (75 pb^{−1}) for conclusion!